### v1

# Adversarial Attacks on Data Attribution

Abstract
Data attribution aims to quantify the contribution of individual training data points to the outputs of an AI model, which has been used to measure the value of training data and compensate data providers. Given the impact on financial decisions and compensation mechanisms, a critical question arises concerning the adversarial robustness of data attribution methods. However, there has been little to no systematic research addressing this issue. In this work, we aim to bridge this gap by detailing a threat model with clear assumptions about the adversary’s goal and capabilities, and by proposing principled adversarial attack methods on data attribution. We present two such methods, Shadow Attack and Outlier Attack, both of which generate manipulated datasets to adversarially inflate the compensation. The Shadow Attack leverages knowledge about the data distribution in the AI applications, and derives adversarial perturbations through “shadow training”, a technique commonly used in membership inference attacks. In contrast, the Outlier Attack does not assume any knowledge about the data distribution and relies solely on black-box queries to the target model’s predictions. It exploits an inductive bias present in many data attribution methods—outlier data points are more likely to be influential—and employs adversarial examples to generate manipulated datasets. Empirically, in image classification and text generation tasks, the Shadow Attack can inflate the data-attribution-based compensation by at least 200%, while the Outlier Attack achieves compensation inflation ranging from 185% to as much as 643%.
1 Introduction
Data attribution aims to quantify the contribution of individual training data points to the outputs of an Artificial Intelligence (AI) model (Koh & Liang, 2017). A key application of data attribution is to measure the value of training data in AI systems, enabling appropriate compensation for data providers (Ghorbani & Zou, 2019; Jia et al., 2019). With the rapid advancement of generative AI, these methods have gained increased relevance, particularly in addressing copyright concerns. Recent studies have explored economic frameworks using data attribution for copyright compensation, showing promising preliminary results (Deng & Ma, 2023; Wang et al., 2024).
Given the significant potential of data attribution methods for data valuation and compensation, an important question arises regarding their adversarial robustness. As these methods influence financial decisions and compensation mechanisms, they may attract malicious actors seeking to manipulate the system for personal gain. This underscores the need to investigate whether data attribution methods can be manipulated and exploited, as their vulnerabilities could lead to unfair compensation, undermining the trustworthiness of the solutions built on top of these methods.
However, adversarial attacks on data attribution methods have received little to no exploration in prior literature. Along with the proposal of a data-attribution-based economic solution for copyright compensation, Deng & Ma (2023) briefly experimented with a few heuristic approaches (e.g., duplicating data samples) for attacking data attribution methods. A systematic study that clearly defines the threat model and develops principled adversarial attack methods has yet to be conducted.
This work presents the first comprehensive study that aims to fill in this gap. We first outline the threat model by detailing the data compensation workflow and precisely specifying the assumptions we made about the adversary’s objective and capabilities. One key assumption is that the data contribution is periodic, and there is certain persistence across consecutive iterations of data attribution, which is the source of knowledge that the adversary could exploit. Motivating examples are provided to justify this assumption in practice. We also assume that the adversary may either have access to the distribution of the data used by the target AI model, or can get black-box queries of the target model’s predictions. Both assumptions are commonly seen in the AI security literature (Shokri et al., 2017; Chen et al., 2017).
Subsequently, we propose two adversarial attack strategies, Shadow Attack and Outlier Attack, respectively relying on different assumptions about the adversary’s capabilities. The Shadow Attack relies on the access to data distribution, and employs the “shadow training” technique commonly used in membership inference attacks (Shokri et al., 2017) to train “shadow models” that imitate the target model. The adversary can then directly perturb their dataset to achieve a higher compensation rate on these shadow models. The Outlier Attack, instead, does not assume knowledge about the data distribution but only relies on black-box queries of the target model’s predictions. The key idea behind this method lies in an inductive bias of many data attribution methods—outlier data points are more likely to be more influential. The proposed Outlier Attack utilizes adversarial examples (Goodfellow et al., 2015; Chen et al., 2017) to generate realistic outliers in a black-box fashion.
We conduct extensive experiments, including both image classification and text generation settings, to demonstrate the effectiveness of the proposed attack methods. Our results show that the Shadow Attack can inflate the adversary’s compensation to at least 200% and up to , while the Outlier Attack can inflate the adversary’s compensation to at least 185% and up to 456%. These results are achieved by adding imperceptible perturbations to real-world data features.
Overall, our study reveals a critical practical challenge in deploying data attribution methods for data valuation and compensation, i.e., the adversarial vulnerability of these methods. Moreover, the design of the proposed adversarial attack methods, especially the Outlier Attack that exploits a common inductive bias of data attribution methods, offers deeper insights into these vulnerabilities. These findings provide valuable directions for future research aimed at enhancing the robustness of data attribution methods.
2 Related Work
Data attribution for data valuation and compensation.
Data attribution methods have been widely used for quantifying the value of training data in AI applications and compensating data providers (Ghorbani & Zou, 2019; Jia et al., 2019; Yoon et al., 2020; Kwon & Zou, 2022; Feldman & Zhang, 2020; Xu et al., 2021; Lin et al., 2022; Kwon & Zou, 2023; Just et al., 2023; Wang & Jia, 2023; Deng & Ma, 2023; Wang et al., 2024). With the rapid advancement of generative AI, these methods have gained increasing relevance due to the growing concerns around copyright. Recent studies have proposed economic solutions using data attribution for copyright compensation, yielding promising preliminary results (Deng & Ma, 2023; Wang et al., 2024). Given the significant potential of data attribution methods for data valuation and compensation, a critical question arises regarding their adversarial robustness. Aside from one earlier exploration using heuristic approaches to attack data attribution methods (Deng & Ma, 2023), no systematic study has addressed this issue. This work presents the first comprehensive study that outlines a detailed threat model and proposes principled and effective approaches for adversarial attacks on data attribution methods.
Membership inference attack.
Membership inference attacks aim to infer whether a specific data point was used during the training of a machine learning model, typically without knowing the actual training dataset or having white-box access to the model (Shokri et al., 2017). The general strategy involves adversaries leveraging information such as model architecture, training data distribution, or black-box model predictions (Shokri et al., 2017; Song & Mittal, 2021). We refer the readers to Hu et al. (2022) for a detailed survey on this topic. One of the proposed attack methods, Shadow Attack, is inspired by the “shadow training” technique (Shokri et al., 2017) commonly used in membership inference attacks, where the adversary draws “shadow samples” following the same distribution as the actual training dataset used by the target model to be attacked, and trains “shadow models” based on the shadow samples to imitate the target model.
Adversarial example.
Adversarial example (Goodfellow et al., 2015) is a well-known phenomenon where small, often imperceptible perturbations to input features can significantly alter the predictions of machine learning models, particularly deep neural networks. These perturbations can be generated through black-box queries to the model predictions (Chen et al., 2017; Ilyas et al., 2018; Guo et al., 2019). For a comprehensive review of adversarial examples, see the surveys by Yuan et al. (2019) and Chakraborty et al. (2021). In the proposed Outlier Attack, we employ black-box adversarial attack methods for generating adversarial examples to generate realistic outliers relative to the training dataset of the target model, without needing access to the training dataset or the model details.
3 The Threat Model
3.1 The Data Compensation Scenario
We consider a scenario where there is an AI Developer, and a (potentially large) set of Data Providers. The Data Providers supply training data for the AI Developer to develop an AI model. In return, the Data Providers are compensated based on their data’s contribution to the model, as measured by a specific data attribution method.
Periodic data contribution.
We assume that the Data Providers contribute data to the AI Developer periodically, a common practice in many AI applications. For example, large language models need periodic updates to stay aligned with the latest factual knowledge about the world (Zhang et al., 2023); recommender systems must adapt to evolving user preferences (Zhang et al., 2020); quantitative trading firms rely on up-to-date information to power their predictive models [^1] See, for example, the Bloomberg market data feed: https://www.bloomberg.com/professional/products/data/enterprise-catalog/market/.; and generative models for music or art benefit from fresh, innovative works by artists to diversify their creative outputs (Smith et al., 2024).
However, such periodic data contribution introduces risks: a malicious Data Provider (referred to as an Adversary thereafter) could exploit information from previous iterations to adversarially manipulate their future data contribution, potentially inflating their compensation unfairly.
To formalize this scenario, without loss of generality, we consider two consecutive iterations of data contribution, denoted as time steps $t=0$ and $t=1$. At $t=0$, there is no Adversary and the training dataset consists solely of contributions from benign Data Providers. This dataset is represented as $Z_{0}\in\mathcal{Z}$, where $\mathcal{Z}$ is the set of all possible training datasets. At $t=1$, the training dataset is $Z_{1}=Z_{1}^{b}\cup Z_{1}^{a}$, where $Z_{1}^{b}\in\mathcal{Z}$ represents the training data provided by benign Data Providers, while $Z_{1}^{a}\in\mathcal{Z}$ is the set provided by the Adversary. We also assume that $Z_{1}^{b}\ \cap Z_{1}^{a}=\varnothing$, meaning there is no overlap between the two datasets.
For any dataset $Z\in\mathcal{Z}$, each data point $z\in Z$ is represented as a pair $z=(x,y)$, where $x\in\mathcal{X}$ is the input feature and $y\in\mathcal{Y}$ is the prediction target. Here, $\mathcal{X}$ and $\mathcal{Y}$ refer to the feature space and target space, respectively.
AI training and data attribution.
Let $\mathcal{M}$ represent the AI model space, and let $\mathcal{T}\colon\mathcal{Z}\to\mathcal{M}$ denote a training algorithm, mapping a training dataset $Z\in\mathcal{Z}$ to a model $\mathcal{T}(Z)\in\mathcal{M}$.
In data attribution, we aim to understand how individual data points from a training dataset $Z\in\mathcal{Z}$ contribute to the model output on a target (validation) data point from a validation dataset $V$. Given $Z$ and $V$, a data attribution method derives a contribution function $\tau\colon Z\times V\to\mathbb{R}$ that assigns a real value $\tau(z,v)$ to each training data point $z\in Z$ for a given validation data point $v\in V$. Denote the set of all such function $\tau$’s as $\mathcal{C}$, and the set of all possible validation sets as $\mathcal{V}$. Therefore, a data attribution method can be formalized as a function $\mathcal{A}\colon\mathcal{Z}\times\mathcal{M}\times\mathcal{V}\to\mathcal{C}$.
Compensation mechanism.
In practice, the contribution function $\tau$ derived from data attribution methods may not reliably measure the contributions of all training data points due to the inherent randomness in AI model training and the need for efficiency (Wang & Jia, 2023; Nguyen et al., 2024). Specifically, measurements of data points with smaller contributions are often less reliable than those of the most influential contributors. Following Deng & Ma (2023), we consider a compensation mechanism where only the top-$k$ influential training data points for each validation data point $v\in V$ receive a fixed amount of compensation.
3.2 The Adversary
We now discuss the objective and capabilities of the Adversary under the data compensation scenario outlined above.
The objective of the Adversary.
Let $V_{1}$ be the validation dataset used at step $t=1$. The objective of the Adversary is to construct a dataset $Z_{1}^{a}$ that maximizes the compensation share received by the Adversary, which is defined as
$c(Z_{1}^{a})=\frac{1}{k|V_{1}|}\sum_{z\in Z_{1}^{a}}\sum_{v\in V_{1}}\mathbf{1}[\tau_{1}(z,v)\in\operatorname{Top}_{k}\left(\{\tau_{1}(z^{\prime},v)\mid z^{\prime}\in Z_{1}\}\right)],$ (1)
where $\tau_{1}=\mathcal{A}(Z_{1},\mathcal{T}(Z_{1}),V_{1})$ is the contribution function at $t=1$; $\operatorname{Top}_{k}(\cdot)$ extracts the top-$k$ elements from a finite set of real numbers; and $\mathbf{1}[\cdot]$ is the indicator function.
The capabilities of the Adversary.
We first outline the limitations imposed on the Adversary. In realistic scenarios, the Adversary does not have access to any of the following:
  • the exact training datasets $Z_{0}$ and $Z_{1}^{b}$;
  • the exact validation datasets $V_{0}$ and $V_{1}$;
  • white-box access to the trained models $\mathcal{T}(Z_{0})$ and $\mathcal{T}(Z_{1})$;
  • the contribution functions $\mathcal{A}(Z_{0},\mathcal{T}(Z_{0}),V_{0})$ and $\mathcal{A}(Z_{1},\mathcal{T}(Z_{1}),V_{1})$.
Then we make the following assumptions to characterize the capabilities of the Adversary.
Assumption 1 (Persistence)
Assume that $Z_{0}\subseteq Z_{1}^{b}$ and $|Z_{0}|/|Z_{1}^{b}|$ is close to 1. Additionally, assume that $V_{0}$ and $V_{1}$ are independently sampled from the same distribution.
Assumption 2 (Access to Data Distribution and Training Algorithm)
Assume that the Adversary has access to the distributions of $Z_{0}$ and $V_{0}$. Additionally, assume that the Adversary has knowledge about the training algorithm $\mathcal{T}$ (but not the model $\mathcal{T}(Z_{0})$).
Assumption 3 (Black-box Access to Model)
Assume that the Adversary has access to black-box query to the model $\mathcal{T}(Z_{0})$, allowing the Adversary to query the model predictions on any input feature $x\in\mathcal{X}$.
Intuitively, 1 assumes persistence between the time steps $t=0$ and $t=1$, so that information gained from $t=0$ can inform the attack at $t=1$. In 2 and 3, it is worth noting that the Adversary’s access to information is limited to what is available at $t=0$ only.
In practice, 1 is realistic in many real-world applications. For example, in applications that have periodic model updates, such as large language models or recommender systems, a substantial portion of the training data often remains consistent across consecutive iterations, with only incremental changes. Moreover, 2 is a common assumption in the membership inference attack literature (Shokri et al., 2017), while 3 reflects a common setup for generating adversarial examples against neural network models (Chakraborty et al., 2021).
The two proposed attack methods in Section 4 and Section 5 respectively rely on 2 and 3, but do not depend on both simultaneously.
The action space of the Adversary.
Finally, we define the action space available to the Adversary. We assume that the Adversary is restricted to making only small perturbations to an existing set of real data points in order to construct the adversarial dataset $Z_{1}^{a}$. This constraint implies that the Adversary cannot introduce entirely synthetic or arbitrary data, but rather, can modify real data points in a subtle manner, which reflects a realistic adversarial scenario where overly large or unnatural alterations would be easily detectable.
4 Shadow Attack
In this section, we introduce the proposed Shadow Attack, which leverages 1 and 2, and exploits the knowledge about the data distribution at $t=0$ to perform attacks.
At a high level, the Adversary first performs a shadow training process, where models are trained on data drawn from a distribution similar to that of the training dataset $Z_{0}$, allowing the Adversary to approximate the target model $\mathcal{T}(Z_{0})$. The Adversary then applies adversarial perturbations to the data points they plan to contribute to the AI Developer at $t=1$. The adversarial perturbations are derived by maximizing the data attribution values on the models obtained through shadow training. See Figure 1 for an illustration.
$M^{(i)}$
$Z\cup Z^{(i)}$
$c_{s}^{\prime}(Z)$
$\tau^{(i)}$
Gradient Ascent
$+\nabla c_{s}^{\prime}$
$Z_{1}^{a}=$
$Z^{(i)}=$
$Z=$
Sampled from $\mathcal{Z}$
$\mathcal{T}$
If converge
Update
Figure 1: An illustration of the Shadow Attack method. Shadow training datasets $Z^{(i)}$’s are sampled to estimate the compensation share of a set of data points $Z$ if it were contributed to the AI Developer, which can be leveraged to perturb data points in $Z$ to get a higher compensation share.
4.1 Shadow Training
Given a dataset $Z\in\mathcal{Z}$ (that the Adversary may eventually contribute to the AI Developer as $Z_{1}^{a}$), the goal of the shadow training process is to estimate the data attribution values of the elements in $Z$ as if this dataset were contributed to the AI Developer at $t=1$.
We first sample $m$ shadow training datasets, denoted as $Z^{(1)},Z^{(2)},\ldots,Z^{(m)}$. These datasets are independent of the actual training dataset used by the AI Developer but follow the same distribution. For each shadow training dataset $Z^{(i)},i=1,\ldots,m$, we train a corresponding shadow model $M^{(i)}\in\mathcal{M}$. The shadow model $M^{(i)}$ is trained on $Z^{(i)}\cup Z$ using the same training algorithm $\mathcal{T}$ as by the AI Developer. [^2] In our experiment in Section 6.2, we demonstrate that the Shadow Attack remains effective when the shadow models have a slightly different architecture compared to the target model. i.e., $M^{(i)}=\mathcal{T}(Z^{(i)}\cup Z)$.
In order to estimate the data attribution values, we further sample a shadow validation dataset $V^{(0)}$ following the same distribution as $V_{0}$. We can estimate a shadow contribution function using each shadow training dataset $Z^{(i)}\cup Z$, the corresponding shadow model $M^{(i)}$, and the shadow validation dataset $V^{(0)}$, resulting in $\tau^{(i)}=\mathcal{A}(Z^{(i)}\cup Z,M^{(i)},V^{(0)})$. Similar to Eq. 1, we consider the following shadow compensation share for the dataset of interest $Z$:
$c_{s}(Z)=\frac{1}{mk|V^{(0)}|}\sum_{i=1}^{m}\sum_{z\in Z}\sum_{v\in V^{(0)}}\mathbf{1}\left[\tau^{(i)}(z,v)\in\operatorname{Top}_{k}(\{\tau^{(i)}(z^{\prime},v)\mid z^{\prime}\in Z^{(i)}\cup Z\})\right].$ (2)
4.2 Adversarial Manipulation by Maximizing Shadow Compensation Rate
Given the shadow compensation rate, a natural idea is to apply adversarial perturbations to $Z$ by solving the following optimization problem
$Z_{1}^{a}=\operatorname*{arg\,max}_{Z^{\prime}\in\mathcal{N}(Z)}c_{s}(Z^{\prime}),$ (3)
where $\mathcal{N}(Z)$ specifies the space of datasets with undetectable perturbations.
In practice, however, solving this optimization problem has two technical challenges. Firstly, the objective $c_{s}(\cdot)$ is discrete and can be difficult to optimize. Secondly, when doing iterative-style optimization, such as gradient ascent, $\tau^{(i)}$’s need to be updated and re-evaluated at every step as it depends on $Z$ and hence $M^{(i)}=T(Z^{(i)}\cup Z)$. This leads to two problems: On the one hand, updating $M^{(i)}$’s requires retraining, which is computationally heavy; on the other hand, many data attribution methods are computationally expensive, even with all the data and (retrained) models available. Hence, solving the problem in Eq. 3 requires repeatedly retraining and running data attributions on perturbed datasets, which may be infeasible for many training algorithms and data attribution methods.
To address the first challenge, we replace $c_{s}(\cdot)$ with the following surrogate objective
$c^{\prime}_{s}(Z)=\frac{1}{mk|V^{(0)}|}\sum_{i=1}^{m}\sum_{z\in Z}\sum_{v\in V^{(0)}}\tau^{(i)}(z,v),$ (4)
which aims to directly maximize the contribution values of the data points in $Z$ on the shadow validation set $V^{(0)}$.
To address the second challenge, we first approximate the retrained models by the initial models trained with the original $Z$ (before any gradient ascent steps) for computational efficiency, and we further adopt Grad-Dot (Charpiat et al., 2019), one of the most efficient data attribution methods when evaluating the contribution function. [^3] Empirically, this works well even when the actual data attribution method used by the AI Developer is more advanced ones, as shown in Section 6.2. Using this method, the contribution value $\tau^{(i)}(z,v)$ for any training data point $z$ and validation data point $v$ equals the dot product between the gradients of the loss function on $z$ and that on $v$, evaluated with model $M^{(i)}$.
With such simplification, the adversarial perturbation can be derived by maximizing Eq. 4, which can be solved efficiently through gradient ascent. Note that since Eq. 4 is a constrained optimization problem, when $\mathcal{N}(Z)$ is bounded, gradient ascent is guaranteed to converge to some local optimums. In practice, we perform a fixed number of iterations and carefully control the overall perturbation budgets when solving Eq. 4. The computation cost for each iteration is approximately the same as one forward and one backward pass on each of the $m$ shadow models.
5 Outlier Attack
For large-scale AI applications, such as generative AI services, 2 might be overly strong, as it could be difficult for the Adversary to guess the distribution of the full training data. In this case, however, 3 often holds true.
In this section, we introduce another proposed attack method, Outlier Attack, which only relies on 1 and 3. See Figure 2 for an illustration.
$\mathcal{T}(Z_{0})$
$Z_{1}^{a}=$
$\ell(Z)$
Black-Box Attack
$Z=$
Collect from $\mathcal{Z}$
Prediction
If increase
enough
$Z$
Black-Box Query
Update
Figure 2: An illustration of the Outlier Attack method. Here, $\ell(Z)$ denotes the loss used by the model $\mathcal{T}(Z_{0})$ when evaluated on the dataset $Z$. The data points in $Z$ are perturbed by maximizing the loss $\ell(Z)$ through black-box attack methods designed to generate adversarial examples.
5.1 The Outlier Inductive Bias of Data Attribution
The core idea behind Outlier Attack leverages an inherent inductive bias present in many data attribution methods: outlier data points in the training dataset tend to be more influential. Indeed, one of the very first applications of the Influence Function developed in the statistic literature was to detect outliers in training data (Cook, 1977). Consequently, if the Adversary contributes a set of outlier data points, they are more likely to get higher compensation, especially given that the compensation mechanism focuses on the top influential data points.
However, translating this intuition into a practical adversarial manipulation strategy poses two significant challenges. Firstly, the contributed data points must closely resemble real-world data; otherwise, they could be easily flagged by data quality filters employed by the AI Developer. Secondly, we aim to develop an attack method that does not rely on direct knowledge about the training dataset or its underlying distribution. This makes it challenging to determine whether certain data points truly qualify as outliers relative to the training dataset gathered by the AI Developer.
5.2 Generating Realistic Outliers with Adversarial Attacks
Next, we propose a general strategy to address the two aforementioned challenges. To tackle the first challenge, we propose starting with a set of real-world data and transforming them into outliers through small adversarial perturbations. With careful design, this strategy also addresses the second challenge: if we can perturb a data point such that an AI model has low confidence in correctly predicting its target, then it is highly likely that this data point behaves as an outlier relative to the model’s training dataset. More specifically, we discuss two key design aspects for generating outliers by perturbing real data points.
  1. 1.
    Data component to be perturbed. Recall that a data point $z=(x,y)$ consists of both the input feature $x$ and the prediction target $y$. To obtain an outlier data point from $z$, we perturb the feature $x$ only, without modifying the target $y$. Intuitively, flipping the annotated target $y$ could easily result in an outlier data point through mislabeling, such an outlier typically degrades model performance and is likely to be identified as negatively influential. Moreover, mislabeled data points are easier for the AI Developer to detect. Therefore, it is more effective to perturb only the feature $x$ while keeping the target $y$ unchanged.
  2. 2.
    Objective of perturbation. For each data point $z=(x,y)$, we aim to decrease the model confidence in predicting the annotated target $y$ by perturbing $x$. In most cases, this is equivalent to increasing the loss $\ell$ between the model prediction and the target $y$.
Combining these two design choices, the resulting perturbed data points coincide with what are commonly referred to as adversarial examples in the literature of adversarial attacks against neural network models. Consequently, existing adversarial attack methods for deriving adversarial examples can be leveraged to generate realistic outliers in our problem setup. Notably, this strategy is very general and can be applied to a variety of data modalities and models by leveraging different off-the-shelf black-box adversarial attack methods.
Choices of adversarial attack methods.
We consider two concrete machine learning settings, image classification, and text generation, and discuss the choices of adversarial attack methods. For smaller-scale image classification settings, we leverage Zeroth Order Optimization (ZOO) (Chen et al., 2017) based adversarial attacks, which approximates the gradient ascend on the loss function with respect to the data features using black-box queries to the target model $\mathcal{T}(Z_{0})$. For larger-scale image classification settings, we employ a more advanced black-box adversarial attack method, Simba (Guo et al., 2019), which is computationally more efficient. At a high level, Simba sequentially perturbs each scalar pixel value in the image by trying to perturb in both directions and accept the perturbation once it increases the loss. For the text generation setting, we utilize TextFooler (Jin et al., 2020), a black-box adversarial attack method tailored for text data. In all the methods of choice, the attack only requires the black-box queries to the predictions of the target model.
6 Experiments
In this section, we empirically evaluate the effectiveness of the two proposed methods, Shadow Attack and Outlier Attack.
6.1 Experimental Setup
Tasks, datasets, target models, and data attribution methods.
We consider two machine learning tasks, image classification, and text generation. For image classification, we experiment on MNIST (LeCun, 1998) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ two data attribution methods, Influence Function (Koh & Liang, 2017) and TRAK (Park et al., 2023). For text generation, we conduct experiments on NanoGPT trained on the Shakespeare dataset (Karpathy, 2022), with TRAK as the data attribution method. For the image classification settings, we evaluate both Shadow Attack and Outlier Attack on image classification tasks, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. The experimental settings are summarized in Table 1.
Data contribution workflow.
For the image classification settings, we set $|Z_{0}|=10000,|Z_{1}^{a}|=100$, and $|Z_{1}|=11000$. This setup simulates the following data contribution workflow: At $t=0$ when there is no Adversary, and $10000$ training points are used to train the model $\mathcal{T}(Z_{0})$. At $t=1$, the Adversary contributes $100$ perturbed training points and other Data Providers contribute $900$ new training data points. Together with the previous $10000$ training points, a total of $11000$ training points are used to train the model $\mathcal{T}(Z_{1})$. The text generation setting follows a similar workflow with $|Z_{0}|=4706,|Z_{1}^{a}|=20$, and $|Z_{1}|=6274$.
Evaluation metrics.
We evaluate the proposed attack methods using two metrics. The first is the Compensation Share, as defined in Eq. 1. For each attack method, we calculate the Compensation Share respectively when the original dataset without perturbation is contributed as $Z_{1}^{a}$ (Original) and when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$ (Manipulated). The increase of the Compensation Share after the perturbation (as measured by the Ratio) reflects the effectiveness of the attack.
To gain a more detailed understanding of how the adversarial perturbations affect the top-$k$ influential data points for individual validation data points in $V_{1}$ we introduce the second evaluation metric named Fraction of Change. For each validation data point $v\in V_{1}$, we measure how many data points in $Z_{1}^{a}$ appear in the top-$k$ influential data points for $v$. In comparison to when the original dataset without perturbation is contributed as $Z_{1}^{a}$, we calculate the fraction of validation data points in $V_{1}$ that contain more data points from $Z_{1}^{a}$ in the top-$k$ influential data points when the manipulated dataset after perturbation is contributed as $Z_{1}^{a}$. Similarly, we calculate the fraction of validation points that contain the same number of or fewer data points from $Z_{1}^{a}$ after perturbation. We report the three fractions under the categories More, Tied, and Fewer for the Fraction of Change metric. The higher fraction for the More category indicates that the attack method influences the validation data points more broadly.
For both metrics, we set $k=100$ when counting the top-$k$ influential data points.
Table 1: Summary of experimental settings with various tasks, datasets, target models, and data attribution methods.
Setting Task Dataset Target Model Attribution Method
(a) Image Classification MNIST LR Influence Function
(b) Image Classification MNIST CNN TRAK
(c) Image Classification CIFAR-10 ResNet-18 TRAK
(d) Text Generation Shakespeare NanoGPT TRAK
6.2 Experimental Results: Shadow Attack
The results of the proposed Shadow Attack method are shown in Table 2. We conduct experiments on all three image classification settings outlined in Table 1. For each of the settings, we first consider the case that the shadow models have the same architecture as the target model, as shown in the first three rows of Table 2. Additionally, for the third setting where the target model is ResNet-18, we further consider using ResNet-9 as the shadow model, which simulates scenarios where the Adversary’s knowledge about the training algorithm is limited.
Table 2: Results of the Shadow Attack method. The target models used for evaluation in each setting are listed in Table 1 while the shadow models used in the attack are listed under the column “Shadow Model”. The proportion of the data contributed by the Adversary relative to the full dataset at $t=1$ is shown under the column “$|Z_{1}^{a}|/|Z_{1}|$”. The Original dataset consists of random samples without perturbation, so its Compensation Share aligns closely with $|Z_{1}^{a}|/|Z_{1}|$. Overall, a higher Ratio indicates a more effective attack. A higher fraction under More means that the attack influences the validation data points more broadly.
Setting Shadow Model $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) LR 0.0098 0.0098 0.0477 456.1% 0.955 0.038 0.007
(b) CNN 0.0098 0.0112 0.0467 417.0% 0.781 0.195 0.024
(c) ResNet-18 0.0098 0.0095 0.0213 217.3% 0.655 0.259 0.086
(c) ResNet-9 0.0098 0.0095 0.0196 206.3% 0.622 0.310 0.068
The Shadow Attack method is highly effective in increasing the Compensation Share across all the settings. The Ratio of the Manipulated to the Original ranges from $206.3\%$ to $456.1\%,$ representing a substantial increase. Notably, the last row corresponds to the setup where the shadow models have the ResNet-9 architecture while the target model is a ResNet-18, and the Shadow Attack remains significantly effective ($206.3\%,$) although being slightly worse than the case where shadow model and target model shares the same architecture (the third row, $217.3\%.$)
The Shadow Attack is also uniformly effective across a wide range of validation data points, as measured by the metrics of Fraction of Change. The attack is able to increase the number of top-$k$ influential points from $Z_{1}^{a}$ for more than $60\%$ of the validation points. In contrast, only less than $9\%$ of the validation data points have fewer top-$k$ influential points from $Z_{1}^{a}$ after the attack.
6.3 Experimental Results: Outlier Attack
The results of the proposed Outlier Attack method are presented in Table 3, where we include all four settings summarized in Table 1. The black-box attack method for generating the adversarial examples is chosen according to the discussion in Section 5.2.
On experimental settings (a) and (b), the Outlier Attack performs even better than the Shadow Attack in terms of both Compensation Share and Fraction of Change, achieving a higher Ratio and a larger fraction under More. Across all four settings, the Ratio ranges from $185.2\%$ to $643.9\%,$ demonstrating the exceptional effectiveness of the Outlier Attack. The results from setting (d), the text generation setting, further highlight the generalizability of the proposed method.
Table 3: Results of the Outlier Attack method. The black-box adversarial attack methods for generating adversarial examples are listed under the column “Attack Method”. See Table 2 for more context.
Setting Attack Method $|Z_{1}^{a}|/|Z_{1}|$ Compensation Share Fraction of Change
Original Manipulated Ratio More Tied Fewer
(a) ZOO 0.0098 0.0098 0.0631 643.9% 0.980 0.017 0.003
(b) Simba 0.0098 0.0112 0.0668 596.4% 0.799 0.173 0.028
(c) Simba 0.0098 0.0095 0.0176 185.2% 0.562 0.354 0.084
(d) TextFooler 0.0031 0.0035 0.0092 262.9% 0.392 0.461 0.147
6.4 Baseline Reference
To better understand the significance of the results in Table 2 and Table 3, we present experimental results for a baseline random perturbation method in Table 4. We conduct experiments on the three settings of image classification, where the baseline method applies pixel-wise random perturbation to the images, with the perturbation budget matched to that of the proposed attack methods. As shown in the results, the Compensation Share of the dataset with Random Perturbation is nearly identical to that of the Original. The fraction under More is also close to that under Fewer. These results further emphasize that the effectiveness of the proposed attack methods comes from the careful design of adversarial perturbation.
Table 4: Results of the random perturbation baseline. See Table 2 for more context.
Setting Compensation Share Fraction of Change
Original Random Perturbation More Tied Fewer
(a) 0.0098 0.0097 0.203 0.586 0.211
(b) 0.0112 0.0117 0.385 0.326 0.289
(c) 0.0095 0.0125 0.367 0.430 0.203
6.5 Visualization of the Perturbation
Finally, we visualize two examples of the adversarially perturbed images in Figure 3 and Figure 4, respectively from MNIST and CIFAR-10. As can be seen, the perturbations are barely visible to human eyes. However, these perturbations make the images significantly more influential for a large number of validation data points.
(a) Original image. Influential for 00 validation data points.
(b) Image after Shadow Attack. Influential for $75$ validation data points.
(c) Image after Outlier Attack. Influential for $105$ validation data points.
Figure 3: Visualization of an image from MNIST, before and after manipulation.
(a) Original image. Influential for $1$ validation data points.
(b) Image after Shadow Attack. Influential for $38$ validation data points.
(c) Image after Outlier Attack. Influential for $29$ validation data points.
Figure 4: Visualization of an image from CIFAR-10, before and after manipulation.
7 Conclusion
In conclusion, this work addresses a significant gap in the current understanding of the adversarial robustness of data attribution methods, which are increasingly influential in data valuation and compensation applications. By introducing a well-defined threat model, we have proposed two novel adversarial attack strategies, Shadow Attack and Outlier Attack, which are designed to manipulate data attribution and inflate compensation. The Shadow Attack utilizes knowledge of the underlying data distribution, while the Outlier Attack operates without such knowledge, relying on black-box queries. Empirical results from image classification and text generation tasks demonstrate the effectiveness of these attacks, with compensation inflation ranging from 185% to 643%. These findings underscore the need for more robust data attribution methods to guard against adversarial exploitation.